OpenStudy (anonymous):

Simplify: (5+3√2) ( 5-3√2)

OpenStudy (kropot72):

This looks like the factorization of the difference of two squares: \[a ^{2}-b ^{2}=(a+b)(a-b)\] So we can find the solution by first doing the following: \[(5+3\sqrt{2})(5-3\sqrt{2})=5^{2}-(3\sqrt{2})^{2}\] Can you simplify \[5^{2}-(3\sqrt{2})^{2}=\ ?\]

OpenStudy (anonymous):

19? @kropot72

OpenStudy (kropot72):

Not really. \[5^{2}=25\] and \[(3\sqrt{2})^{2}=18\] So we get \[25-18=\ ?\]

OpenStudy (anonymous):

7 @kropot72

OpenStudy (nirmalnema):

yup 7 is correct

OpenStudy (anonymous):

can you help me @nirmalnema

OpenStudy (nirmalnema):

ya sure

OpenStudy (anonymous):

(5x-1)^2=24

OpenStudy (nirmalnema):

a=5x b=-1 (a-b)^2=a^2 - 2ab + b^2 (5x-1)^2= 25x^2-10x+1=24 25x^2-10x-23=0

OpenStudy (nirmalnema):

use the formula to solve the quadratic \[(-b(\pm)\sqrt{b^2-4ac})/2a\]

OpenStudy (nirmalnema):

a=25 b=-10 c=-23

OpenStudy (anonymous):

can u solve it all the way? @nirmalnema

OpenStudy (mrnood):

@cutegirl I'm sure he can. Can YOU?

OpenStudy (nirmalnema):

yaa..1st you try by yourself if you are not able to do then i will surely help u

OpenStudy (anonymous):

2√6/5+1/5 ??

OpenStudy (nirmalnema):

you will get 2 ans 1 one by taking + sign and other by taking - sign

OpenStudy (anonymous):

1/5 -2√6/5

OpenStudy (anonymous):

can you help me with this one? (√7+√x) /(√7-√x) @Nicole143